### v1

# On computational complexity and average-case hardness of shallow-depth boson sampling

Abstract
Boson sampling, a computational task believed to be classically hard to simulate, is expected to hold promise for demonstrating quantum computational advantage using near-term quantum devices. However, noise in experimental implementations poses a significant challenge, potentially rendering boson sampling classically simulable and compromising its classical intractability. Numerous studies have proposed classical algorithms under various noise models that can efficiently simulate boson sampling as noise rates increase with circuit depth. To address this issue particularly related to circuit depth, we explore the viability of achieving quantum computational advantage through boson sampling with shallow-depth linear optical circuits. Specifically, as the average-case hardness of estimating output probabilities of boson sampling is a crucial ingredient in demonstrating its classical intractability, we make progress on establishing the average-case hardness confined to logarithmic-depth regimes. We also obtain the average-case hardness for logarithmic-depth Fock-state boson sampling subject to lossy environments and for the logarithmic-depth Gaussian boson sampling. By providing complexity-theoretical backgrounds for the classical simulation hardness of logarithmic-depth boson sampling, we expect that our findings will mark a crucial step towards a more noise-tolerant demonstration of quantum advantage with shallow-depth boson sampling.
preprint: APS/123-QED
I Introduction
Boson sampling is a computational task that is complexity-theoretically proven to be hard to classically simulate under plausible assumptions [1, 2, 3]. Accordingly, boson sampling has gathered significant attention, as it would possibly play a key role in the experimental demonstration of quantum computational advantage using near-term quantum devices. However, the implementation of boson sampling in experimental settings with near-term quantum devices is inevitably subject to various sources of noise [4, 5, 6, 7]. The problem is that those noises would possibly rule out the classical intractability of boson sampling, and thus potentially hinder the experimental demonstration of quantum advantage with boson sampling. Indeed, both for finite-size near-term experiments and asymptotic limits as system size scales, numerous studies [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] have proposed efficient classical simulation algorithms of boson sampling under various noise models, such as photon loss, partial distinguishability, gaussian noise, etc. Their results indicate that as the noise rate of boson sampler increases, it eventually renders such a noisy sampler classically simulable. Moreover, as the noise is typically accumulated with each circuit depth, the quantum signal for classical intractability exhibits exponential decay with increasing circuit depth. Hence, circuits with polynomially increasing depth with system size would suffer from significantly enlarged noise rates, posing substantial challenges to achieving quantum advantage in such settings.
[…]
In this work, we investigate the classical simulation hardness of boson sampling in shallow linear optical circuits. Specifically, as the average-case #P-hardness of estimating output probabilities of boson sampling is a crucial ingredient to demonstrate the classical intractability of boson sampling, we make progress on establishing the average-case #P-hardness confined in shallow-depth regimes. Similarly, we obtain the average-case hardness result in the shallow-depth regime for the Gaussian boson sampling scheme. Finally, since noise is our main motivation for investigating shallow-depth boson sampling, we generalize our average-case hardness result to noisy boson sampling subject to a photon loss channel.
To avoid confusion, we note that the allowed imprecision level of our average-case #P-hardness result is not sufficient to fully demonstrate the classical intractability of boson sampling in shallow-depth regimes. However, to the best of our knowledge, the complexity-theoretical analysis on the average-case hardness of shallow-depth boson sampling has not yet been investigated. Hence, we believe that our hardness result in shallow-depth regimes will provide a first step toward a stronger hardness result and, ultimately, toward the full demonstration of the classical intractability of shallow-depth boson sampling.
I.1 Outlines: average-case hardness of shallow-depth boson sampling
We set our goal as proving the hardness of classical simulation of boson sampling in the shallow-depth regime, specifically for approximate simulation within total variation distance error. Two key ingredients for the current hardness proof of the approximate simulation of boson sampling are $(i)$ average-case #P-hardness of output probability approximation up to sufficiently large additive imprecision $\epsilon$, and $(ii)$ hiding property. Informally, average-case hardness means that approximating the output probability of boson sampling with high probability over randomly chosen circuits (i.e., on average over circuits) is #P-hard. Here, by choosing random circuit instances that have the hiding property (i.e., symmetric over outcomes), one can reduce the average-case instances for the hardness from circuit instances to outcome instances, which is a crucial step to prove the hardness of approximate simulation within total variation distance error (See Appendix A for more details).
Most of the current theoretical foundations of the average-case hardness of boson sampling rely on global Haar random unitary circuits [1, 2, 3, 20, 21, 22], as they almost satisfy the two conditions described above. Namely, the outcome instances can be effectively hidden by global Haar random unitaries, and approximating the output probability within sufficiently large $\epsilon$ on average over global Haar random circuit instances is #P-hard under some conjectures. However, the problem is that implementing global Haar random unitary requires at least polynomial circuit depth (e.g., see [23, 24]), and thus not implementable in sub-polynomial circuit depths. Accordingly, the hardness results built upon global Haar random unitaries cannot be directly applied to the shallow-depth boson sampling we are interested in, necessitating a different approach from them.
[…]
To deal with those problems we take the following approach: first, we consider shallow linear optical circuit architectures composed of geometrically non-local gates. In fact, implementation of non-local gates is promising for near-term experimental settings; for example, experiments of linear optical systems based on trapped ions [30, 31] and photonic architecture [6] implemented long-range interactions. Also, since we cannot implement global Haar random unitary within shallow-depth regime, we instead employ local random circuit ensemble for random circuit instances in shallow circuit architecture, inspired by the hardness results of random circuit sampling [32, 33, 21, 34, 35]. Here, local random distribution in this context means that each gate composing the circuit is independently chosen Haar random gate; we note that recent experimental setups of boson sampling [4, 5, 6, 7] also follow such circuit distribution, but with geometrically local architectures.
However, local random distribution poses a subsequent challenge toward the average-case hardness of shallow-depth boson sampling, that is, the absence of the hiding property. Since the output symmetry of boson sampling over local random circuit distribution is not evident, the random outcome instances cannot be efficiently hidden by random circuit instances. This means that even if we find the average-case hardness over randomly chosen circuits for a fixed outcome, it still does not lead to the classical simulation hardness grounded in Stockmeyer’s reduction from the average-case approximation over randomly chosen outcomes [1]. To address this issue, we prove the average-case hardness over both the randomly chosen outcome and randomly chosen circuit, by establishing a worst-to-average-case reduction for both outcome and circuit instances. Specifically, our reduction process is composed of two steps: $(i)$ from a given fixed outcome to a randomly chosen collision-free outcome, and $(ii)$ from a worst-case circuit to a randomly chosen circuit over local random circuit distribution.
To sum up, we show the average-case hardness over outcomes and circuit instances for shallow circuit architectures composed of non-local gates and employing the local random circuit ensemble. We informally present here our average-case hardness result of boson sampling in the logarithmic depth regime, for photon number $N$ and mode number $M\propto N^{\gamma}$ with $\gamma\geq 1$.
Theorem 1 (Informal).
There exists a $O(\log N)$-depth linear optical circuit architecture such that approximating output probability of boson sampling within additive error $2^{-O(N^{\gamma+1}(\log N)^{2})}$ with high probability over randomly chosen circuits in the circuit architecture for high probability over randomly chosen collision-free outcomes is #$\rm{P}$-hard under $\rm{BPP}^{\rm{NP}}$ reduction.
Also, since our average-case hardness result considers both the random outcomes and random circuits due to the absence of the hiding property, it is not straightforward to show the classical simulation hardness of shallow-depth boson sampling as in the original boson sampling proposal [1]. Accordingly, we show how our average-case hardness result over both the randomly chosen outcomes and circuits leads to the classical simulation hardness argument. This implies that improving the additive imprecision for our average-case hardness result is the only remaining problem for the fully theoretically guaranteed classical intractability of shallow-depth boson sampling.
Theorem 2 (Informal).
If the allowed additive imprecision for the problem in Theorem 1 to be #$\rm{P}$-hard is improved to $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$, the approximate boson sampling for the shallow-depth circuit in Theorem 1, up to constant total variation distance, is classically hard to simulate.
Now we provide an outline of our results, which is depicted in Fig. 1. We first define in Sec. II a shallow-depth circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ composed of non-local gates, which we will use throughout our results. Next, in Sec. III we prove the worst-case #P-hardness of approximating output probability $p_{\bm{s}}(C)$ of a fixed outcome $\bm{s}$ of boson sampling, for any circuit $C$ in the shallow circuit architecture previously defined. In Sec. IV we prove the average-case #P-hardness of approximating output probability $p_{\bm{s}}(U)$ for randomly chosen outcome $\bm{s}$ and randomly chosen circuit $U$ in the shallow circuit architecture, by establishing worst-to-average-case reduction. We prove in Sec. V how our average-case hardness results over both the random outcomes and random circuits lead to the classical simulation hardness. We also extend our average-case hardness result to the Gaussian boson sampling scheme in Sec. VI, and to the lossy boson sampling subject to photon loss channels in Sec. VII. In Sec. VIII, we conclude with several remarks.
Figure 1: Outlines of our result
II Notations
Let us define the total mode number as $M$, where we set $M$ as a power of 2 for simplicity. We set the output photon number $N$ polynomially related to $M$ as $M=c_{0}N^{\gamma}$, for a constant $c_{0}$ and $\gamma\geq 1$ satisfying $M\geq 2N$. We use the notation $\bm{s}$ as an $M$-dimensional output configuration vector for the collision-free outcome, such that each element $s_{i}$ of $\bm{s}$ denotes photon number in $i$th mode. Namely, $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}s_{i}=N$, so that the number of possible configurations of $\bm{s}$ is $\binom{M}{N}$. We define $p_{\bm{s}}(C)$ as an output probability of a linear optical circuit (unitary) matrix $C$ for the outcome $\bm{s}$ from a predefined input configuration $\bm{t}$. For collision-free input and output, $p_{\bm{s}}(C)$ can be represented as [1]
$p_{\bm{s}}(C)=|\text{Per}(C_{\bm{s},\bm{t}})|^{2}$ (1)
where $C_{\bm{s},\bm{t}}$ is a $N$ by $N$ matrix obtained by taking $s_{i}$ copies of the $i$th row and $t_{j}$ copies of the $j$th column of the matrix $C$.
We note that an $M$-mode linear optical circuit can be represented by an $M$ by $M$ unitary matrix in U$(M)$ which unitarily transforms $M$ mode operators. Specifically, we can represent a single-mode gate (i.e., a phase shifter) as a U$(1)$ matrix to the mode, and a two-mode gate (i.e., a beam splitter) as a U$(2)$ matrix along the modes. Also, the parallel application of gates can be represented as a unitary matrix with a block matrix form, and the serial application of gates can be represented as matrix multiplication of the unitary matrices. Accordingly, throughout this work, we will interchangeably use the terminology ‘(linear optical) circuit’ and ‘(unitary) matrix’.
We first define the linear optical circuit architecture, for a more rigorous analysis of the hardness proof.
Definition 1 (Linear optical circuit architecture).
The linear optical circuit architecture $\mathcal{A}$ is a linear optical circuit with fixed type (i.e., single- or two-mode) and fixed location of gates, where the coefficients of each gate are not specified. If the coefficients of each gate are specified with unitary matrices (in $\rm{U}$$(1)$ or $\rm{U}$$(2)$), then the circuit and the corresponding unitary matrix are specified.
Figure 2: Schematics of the butterfly circuit architectures in Definition 2 and their unitary matrix form, for mode number $M=2^{4}=16$.
For the shallow-depth circuit architecture, specifically in logarithmic depth, we define the shallow linear optical circuit architecture of circuit depth $D=\log M$, using the convention used in [36, 37].
Definition 2.
We define butterfly circuit architecture $\mathcal{B}$ as follows: for each layer $L=1,2,\dots,D=\log M$ of the circuit architecture, allocate two-mode gate between mode number $2^{L}(j-1)+k$ and $2^{L}(j-1)+k+2^{L-1}$, for all $j=1,2,\dots,2^{D-L}$ and $k=1,2,\dots,2^{L-1}$. Also, we define inverse butterfly circuit architecture $\mathcal{B}^{*}$ as a butterfly circuit architecture with the inverse sequence of gate application along the depth.
We illustrate in Fig. 2 the circuit architecture $\mathcal{B}$ and $\mathcal{B}^{*}$, and the form of their corresponding unitary matrix. Next, we define the Kaleidoscope circuit architecture proposed in [36], using the butterfly circuit architecture defined above.
Definition 3.
We define Kaleidoscope circuit architecture $\mathcal{B}\mathcal{B}^{*}$ as a serial application of $\mathcal{B}$ over $\mathcal{B}^{*}$. We also define $q$-Kaleidoscope circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ as a repeat of the Kaleidoscope circuit architecture, with repetition number $q\in\mathbb{N}$.
Here, the circuit depth of $q$-Kaleidoscope architecture is $D=2q\log M$, which is indeed a logarithmic depth in $N$ for $q=O(1)$. Throughout this paper, we will focus on the $q$-Kaleidoscope circuit architecture with $q=O(1)$ to demonstrate the hardness results for shallow-depth circuits. One motivation for employing this linear optical circuit architecture is that it enjoys a useful property that is crucial for our analysis; Ref. [36] shows that for $M$ a power of $2$, any $M$ mode permutation circuit can be implemented within $\mathcal{B}\mathcal{B}^{*}$.
Lemma 1 (Dao et al [36]).
Let $\bm{P}$ be an arbitrary $M\times M$ permutation matrix with $M$ a power of 2. Then $\bm{P}$ can be efficiently implemented in $\mathcal{B}\mathcal{B}^{*}$ using two-mode permutation gates, i.e., $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$.
III Worst-case hardness of output probability estimation
In this section, we find the worst-case #P-hardness of output probability estimation of shallow-depth linear optical circuits in $\mathcal{B}\mathcal{B}^{*}$, for a fixed input and output $\bm{s}$ within a certain additive imprecision. Our worst-case hardness result for the shallow circuit architecture can be represented as follows.
Theorem 3 (Worst-case hardness).
For $M\geq 2N$, approximating the output probability $p_{\bm{s}}(C)$ to within additive error $2^{-O(N)}$ for any $C$ over linear optical circuit architecture $\mathcal{B}\mathcal{B}^{*}$ is #$\rm{P}$-hard in the worst case.
We briefly sketch the proof of our worst-case hardness result for the shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$. The proof is based on the result by [25], which showed the simulation hardness of exact boson sampling with constant depth linear optical circuits. Specifically, there exist constant-depth linear optical circuits that can simulate an arbitrary given quantum circuit using post-selection. Also, those constant depth circuits can be embedded in our circuit architecture $\mathcal{B}\mathcal{B}^{*}$. Hence, additively approximating the output probability of any quantum circuit can be reduced to additively approximating the output probability of any circuit in $\mathcal{B}\mathcal{B}^{*}$, with imprecision blowup up to the inverse of post-selection probability. Using the fact that the additive approximation of any quantum circuit is #P-hard for certain additive imprecision [34], we can obtain the worst-case hardness of our shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
Proof of Theorem 3.
See Appendix B. ∎
IV Average-case hardness of output probability estimation
In this section, we prove the average-case #P-hardness of approximating output probabilities of shallow-depth boson sampling within a certain additive imprecision. We focus on $q$-Kaleidoscope circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $q=O(1)$, where the gate number $m$ for such architecture is $m=qM\log M$. From the result of Theorem 3, $(\mathcal{B}\mathcal{B}^{*})^{q}$ has the worst-case hardness of output probability approximation, for $q\geq 1$.
[…]
Definition 4 (Random circuit ensemble).
Let $\mathcal{A}$ be the circuit architecture with $m$ number of gates. We define $\mathcal{H}_{\mathcal{A}}$ as the distribution over circuits with architecture $\mathcal{A}$, whose gates are independently distributed local Haar random matrices $\{H_{i}\}_{i=1}^{m}$.
Definition 5 (Random collision-free outcome ensemble).
We define $\,\mathcal{G}_{M,N}$ as the uniform distribution over $\binom{M}{N}$ collision-free outcomes of boson sampling with $M$ modes and $N$ photons. Each outcome $\bm{s}\sim\mathcal{G}_{M,N}$ is an $M$-dimensional output configuration vector for the collision-free outcome, such that $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}s_{i}=N$.
Using the definitions above, we first state our result on the average-case hardness over the outcome and circuit instances, for shallow-depth linear optical circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ with $q=O(1)$.
Theorem 4 (Average-case hardness).
The following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a random circuit $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ and a random outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the output probability $p_{\bm{s}}(U)$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
In the following, we sketch the proof of Theorem 4, by briefly describing the worst-to-average-case reduction process; we leave in Appendix C a detailed proof of Theorem 4. Since our average-case hardness regards both outcomes and circuits, we first describe how to effectively fix the outcome, so that the remaining problem is to establish worst-to-average-case reduction for fixed output probability over random circuit instances. To do so, our strategy is to randomly permute both a given worst-case circuit and a given fixed outcome. That is, we sample random $M$-mode permutation $\bm{P}$ and permute the worst-case circuit $C_{0}$ and the fixed outcome $\bm{s}_{0}$ equally with $\bm{P}$, where the permuted outcome $\bm{s}=\bm{P}\bm{s}_{0}$ now follows the random outcome ensemble $\mathcal{G}_{M,N}$. Then the fixed worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ is equal to $p_{\bm{s}}(\bm{P}C_{0})$, and thus we can obtain the value $p_{\bm{s}_{0}}(C_{0})$ by inferring $p_{\bm{s}}(\bm{P}C_{0})$ via worst-to-average-case reduction over random circuit instances. Here, $p_{\bm{s}}(\bm{P}C_{0})$ now becomes a new worst-case output probability, such that the revised worst-case circuit $C$ is the randomly permuted circuit $\bm{P}C_{0}$, and the revised fixed outcome $\bm{s}$ is the randomly chosen outcome from $\mathcal{G}_{M,N}$.
To establish the worst-to-average-case reduction from the revised worst-case circuit $C$ to the average-case circuit over $\mathcal{H}_{\mathcal{A}}$ for a fixed outcome $\bm{s}$, our strategy is to perturb the circuit from $\mathcal{H}_{\mathcal{A}}$ with the given worst-case circuit $C$ parameterized by a constant $\theta\in[0,1]$. That is, $\theta=0$ corresponds to the average-case distribution $\mathcal{H}_{\mathcal{A}}$ and $\theta=1$ corresponds to the worst-case circuit $C$. Specifically, we choose a perturbation method such that for small enough $\theta$, the success probability of average-case approximation over perturbed random circuits would not have largely deviated from the ideal case ($\theta=0$), and as $\theta$ grows, the perturbed circuit converges to the worst-case circuit $C$. Using such perturbation method and as long as $\theta$ values are small enough, one can obtain the average-case approximate output probability values with high probability over perturbed circuits, parameterized by different values of $\theta$. Also, we can expect that those average-case values contain some information about the worst-case value $p_{\bm{s}}(C)$, depending on the perturbation method and the values of $\theta$. Assuming that worst-case value $p_{\bm{s}}(C)$ can be inferred using the average-case values with small values of $\theta$, one can finally infer the worst-case value $p_{\bm{s}}(C)$, within a certain imprecision determined by the average-case approximation imprecision and the method for the inference.
Therefore, it is crucial to choose a proper perturbation method to establish the worst-to-average-case reduction successfully. Throughout this work, we use the Cayley path for the perturbation, which was employed in Refs. [33, 21, 34] for the hardness proposals of the random circuit sampling.
Definition 6 (Cayley transform [33]).
The Cayley transform of an $n$ by $n$ unitary matrix $H$ parameterized by $\theta\in[0,1]$ is a unitary matrix defined as
$H(\theta)\coloneqq((2-\theta)H+\theta I_{n})(\theta H+(2-\theta)I_{n})^{-1},$ (2)
where $I_{n}$ is the $n$ by $n$ identity matrix. Also, for the diagonalization of the $n$ by $n$ unitary matrix $H=LDL^{{\dagger}}$, with unitary matrix $L$ and diagonal matrix $D=\text{diag}(e^{i\phi_{1}},\dots,e^{i\phi_{n}})$, the equivalent form of the Cayley transform is
$H(\theta)=\frac{1}{q(\theta)}L\;\text{diag}(\left\{p_{j}(\theta)\right\}_{j=1}^{n})\,{\it{L}}^{{\dagger}},$ (3)
where
$q(\theta)=\prod_{j=1}^{n}(1+i\theta e^{i\frac{\phi_{j}}{2}}\sin{\frac{\phi_{j}}{2}}),$ (4)
$p_{j}(\theta)=e^{i\phi_{j}}(1-i\theta e^{-i\frac{\phi_{j}}{2}}\sin{\frac{\phi_{j}}{2}})\prod_{k\in[n]\setminus j}(1+i\theta e^{i\frac{\phi_{k}}{2}}\sin{\frac{\phi_{k}}{2}}).$ ) ) . (5)
Using the Cayley transform defined above, we now define the perturbed random circuit distribution.
Definition 7 (Perturbed random circuit ensemble).
Let $\mathcal{A}$ be the circuit architecture with $m$ number of gates. For the given circuit $C_{0}$ in $\mathcal{A}$ with gates $\{G_{i}\}_{i=1}^{m}$, the circuit $U(\theta)$ is defined with each gate of $C_{0}$ replaced by $G_{i}\rightarrow H_{i}(\theta)G_{i}$, where each $H_{i}(\theta)$ is a Cayley transform of independently distributed local Haar random gate $H_{i}$ $(i\in[m])$ parameterized by $\theta\in[0,1]$. We define $\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ as the distribution for such $U(\theta)$. Here, the distribution of the $U(0)$ is $\mathcal{H}_{\mathcal{A}}$, and $U(1)=C_{0}$.
Before proceeding, we should make sure that the success probability of average-case approximation over circuits is still large enough after the perturbation, to establish the reduction process successfully. This is evident in the case that the total variation distance over circuits induced by the perturbation is small enough, as the success probability over circuits perturbs by, at most, the total variation distance.
In fact, Ref. [33] proved that total variation distance between $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ is small for comparably small perturbation $\theta$.
Lemma 2 (Movassagh [33]).
Let $\mathcal{A}$ be the circuit architecture with $m$ number of gates. For $\theta\ll 1$ and for any circuit $C_{0}$ in $\mathcal{A}$, total variation distance between $\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ and $\mathcal{H}_{\mathcal{A}}$ is $O(m\theta)$.
Therefore, by using small $\theta=O(m^{-1})$, one can upper-bound the total variation distance by an arbitrarily small constant, which implies that the success probability of average-case approximation over circuits also perturbs by at most a small constant.
For the worst-to-average-case reduction, we first sample a random circuit $U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ and worst-case circuit $C_{0}$. Using $U(\theta)$ with the same random seed $\{H_{i}\}_{i=1}^{m}$ but with different values of $\theta$ satisfying $\theta=O(m^{-1})$, we obtain the average-case approximation of the output probability $p_{\bm{s}}(U(\theta))$ for each $\theta$, which may enable us to infer the worst-case output probability value $p_{\bm{s}}(U(1))=p_{\bm{s}}(C)$.
To investigate the feasibility of the inference of the worst-case value, we examine the behavior of the function $p_{\bm{s}}(U(\theta))$ characterized by the parameter $\theta$. Using Definition 6, we find that $p_{\bm{s}}(U(\theta))$ can be represented as a low-degree rational function in $\theta$.
Lemma 3.
Let $\mathcal{A}$ be the $q$-Kaleidoscope circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $m=qM\log M$ number of gates, and $U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ for any $C_{0}$ in $\mathcal{A}$. Then for any outcome $\bm{s}$, the output probability $p_{\bm{s}}(U(\theta))$ can be represented as a degree $(4mN,4mN)$ rational funtion in $\theta$.
Proof.
For given circuit unitary matrix $U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ with $C_{0}$ composed of $\{G_{i}\}_{i=1}^{m}$ gates, one can decompose $U(\theta)$ with $m=qM\log M$ product of unitary matrices, such that each matrix element of $U(\theta)$ can be represented as
$\left[U(\theta)\right]_{j,k}=\sum_{l_{1}=1}^{M}\sum_{l_{2}=1}^{M}\cdots\sum_{l_{m-1}=1}^{M}U_{j,l_{1}}^{(1)}U_{l_{1},l_{2}}^{(2)}\cdots U_{l_{m-1},k}^{(m)},$ (6)
where each $U^{(i)}$ denotes an $M$-dimensional unitary matrix, with a single gate unitary matrix $H_{i}(\theta)G_{i}$ applied to the modes participating in the gate and identity for the rest of the modes. For example, if the $i$th gate $H_{i}(\theta)G_{i}$ is a two-mode gate between the first two modes, $U^{(i)}$ is a block diagonal matrix of $H_{i}(\theta)G_{i}$ and identity matrix, namely, $U^{(i)}=H_{i}(\theta)G_{i}\bigoplus I_{M-2}$.
For circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ which is composed of only two-mode gates, matrix elements of $U^{(i)}$ can be represented as degree $(2,2)$ rational functions in $\theta$, where the common denominator for the elements is given by $q_{i}(\theta)$, defined in Eq. (4) but with index $i$ appended for $i$th random gate $H_{i}(\theta)$. Using reduction to the common denominator for all of the $m$ gates, $\left[U(\theta)\right]_{j,k}$ can be represented as $(2m,2m)$ rational function in $\theta$ with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)$; note that it does not change with the indices $j,k$.
From Eq. (1), the output probability $p_{\bm{s}}(U(\theta))$ has the form of
$\begin{split}p_{\bm{s}}(U(\theta))&=\left|\sum_{\sigma\in S_{N}}\prod_{j=1}^{N}\left[U(\theta)_{\bm{s},\bm{t}}\right]_{\sigma_{j},j}\right|^{2},\end{split}$ (7)
where $\bm{t}$ is an input configuration vector, and $S_{N}$ is $N$-mode permutation group. One can easily check that the common denominator for $\prod_{j=1}^{N}\left[U(\theta)_{\bm{s},\bm{t}}\right]_{\sigma_{j},j}$ is $[\prod_{i=1}^{m}q_{i}(\theta)]^{N}$, and it does not change with permutation $\sigma$. Let us define $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$, which is a degree $4mN$ polynomial in $\theta$. Then $Q(\theta)$ serves as the common denominator for the output probability. Hence, the output probability can be represented as $p_{\bm{s}}(U(\theta))=\frac{P(\theta)}{Q(\theta)}$, with $P(\theta)$ also a degree $4mN$ polynomial in $\theta$. ∎
We are now ready to turn to the proof of Theorem 4, i.e., the average-case hardness of the shallow-depth boson sampling. We prove that for high probability over $\bm{s}\sim\mathcal{G}_{M,N}$, well-approximating output probability $p_{\bm{s}}(U)$ with high probability over $U\sim\mathcal{H}_{\mathcal{A}}$ for the shallow-depth architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ is #P-hard under a $\rm{BPP}^{\rm{NP}}$ reduction.
Proof of Theorem 4.
See Appendix C
V Average-case hardness implies classical simulation hardness
As we have previously discussed, since our average-case hardness result considers both the random outcomes and random circuits, it is not straightforward to show the classical simulation hardness of shallow-depth boson sampling as in the original boson sampling proposal [1]. Therefore, in this section, we provide a self-contained analysis of how our average-case hardness result leads to the classical simulation hardness arguments of shallow-depth boson sampling. Specifically, we show that if the allowed additive error in Theorem 4 for the hardness is improved to a certain imprecision level, an efficient classical algorithm that can approximately simulate the shallow-depth boson sampling is unlikely to exist. This emphasizes that improving the imprecision level of the average-case hardness in Theorem 4 is a crucial step for the classical intractability of shallow-depth boson sampling.
Similarly to Refs. [1, 32], we define an approximate boson sampler as follows.
Definition 8 (Approximate boson sampler).
Approximate boson sampler is a classical randomized algorithm that takes input linear optical circuit $C$ and outputs a sample from the output distribution $\mathcal{D}_{C}^{\prime}$ such that
$||\mathcal{D}_{C}^{\prime}-\mathcal{D}_{C}||\leq\beta$ (8)
where $\mathcal{D}_{C}$ is the ideal output distribution of the circuit $C$ and |||||$\cdot$||| | ⋅ | | represents total variation distance.
Given the total variation distance error, the above approximate sampler can have an arbitrarily large additive error for a fixed output probability. Nevertheless, it still has a comparably small additive error for most of the output probabilities due to Markov’s inequality. Accordingly, finding the average-case solution of the output probability of the ideal sampler over randomly chosen collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, up to a certain additive imprecision, is in complexity class $\rm{BPP}^{\rm{NP}}$ by Stockmeyer’s theorem [38].
Lemma 4 (Average-case approximation [1]).
If there exists an approximate boson sampler $\mathcal{S}$ with total variation distance $\beta$, then for any linear optical circuit $C$, the following problem is in $\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$: find the average-case approximate solution $\tilde{p}_{\bm{s}}(C)$ of $p_{\bm{s}}(C)$, which satisfies
$\Pr_{\bm{s}\sim\mathcal{G}_{M,N}}\left[|\tilde{p}_{\bm{s}}(C)-p_{\bm{s}}(C)|\geq\frac{\kappa}{\binom{M}{N}}\right]\leq\xi,$ (9)
where $\bm{s}$ is over all collision-free outcomes, and $\kappa,\xi>0$ are the fixed error parameters satisfying $\beta=\kappa\xi/{12}$.
We leave the proof of Lemma 4 in Appendix D for a more self-contained analysis. The complexity $\rm{BPP}^{\rm{NP}}$ is known to be inside the finite level of PH, i.e., $\rm{BPP}^{\rm{NP}}\subseteq PH$. Also, by Toda’s theorem [39], PH problems can be solved given the ability to solve any #P problem, i.e., $\rm{BPP}^{\rm{NP}}\subseteq\rm{PH}\subseteq\rm{P}^{\rm{\#P}}$. If finding the average-case solution of output probabilities of sampler $\mathcal{S}$ is #P-hard, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Therefore, if an efficient classical algorithm exists that can simulate $\mathcal{S}$, then $\rm{P}^{\rm{\#P}}\subseteq\rm{BPP}^{\rm{NP}}$ which implies the collapse of PH. Consequently, under the assumption of the non-collapse of the PH, there is no efficient classical algorithm capable of simulating $\mathcal{S}$.
Based on the above arguments, we show that for the case that allowed additive imprecision of Theorem 4 for the hardness can be improved, then it is classically hard to simulate shallow-depth boson sampling within a constant total variation distance.
Theorem 5.
Suppose that the allowed additive imprecision for the problem in Theorem 4 to be #$\rm{P}$-hard can be improved to $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$. Then the efficient classical simulation of approximate boson sampler $\mathcal{S}$ with respect to circuits from the shallow architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ implies the collapse of $\rm{PH}$.
Proof.
We establish a reduction from the problem in Theorem 4 with allowed additive error $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$ to the problem in Lemma 4. Let $\mathcal{O}$ be an oracle that solves the problem in Lemma 4, i.e., on input a circuit $C$ and a randomly chosen collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, $\mathcal{O}$ outputs an estimate of output probability $p_{\bm{s}}(C)$ up to additive imprecision $\kappa\binom{M}{N}^{-1}$, with probability at least $1-\xi$ over outcomes for any circuit. For convenience, we refer to the outcomes which $\mathcal{O}$ estimates with error larger than $\kappa\binom{M}{N}^{-1}$ as bad outcomes, such that the portion of bad outcomes over possible collision-free outcomes is at most $\xi$ for any circuit.
For the randomly chosen circuit input, bad outcomes can vary with the circuit instances, as the sampler $\mathcal{S}$ has the freedom to choose its error distribution according to the input circuit. However, no matter how the bad outcomes vary with circuit instances, $\mathcal{O}$ succeeds at least $1-\frac{\xi}{\eta}$ fraction over circuit instances for at least $1-\eta$ fraction of the outcomes, for any $\eta$ satisfying $\xi<\eta<1$. Otherwise, the failure probability over outcomes and circuits would exceed $\xi$, which contradicts the proposition that the success probability of $\mathcal{O}$ is at least $1-\xi$ over randomly chosen outcomes and circuits. As the above argument holds for any random circuit families, we choose the random circuit distribution as $\mathcal{H}_{\mathcal{A}}$ with the shallow architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$.
To sum up, on input a random circuit $H\sim\mathcal{H}_{\mathcal{A}}$ and a random outcome $\bm{s}\sim\mathcal{G}_{M,N}$, the oracle $\mathcal{O}$ estimates the output probability $p_{\bm{s}}(H)$ up to imprecision $\kappa\binom{M}{N}^{-1}$, with probability at least $1-\frac{\xi}{\eta}$ over the choice of $H$ for at least $1-\eta$ over the choice of $\bm{s}$. Here, the additive imprecision can be bounded as
$\kappa\binom{M}{N}^{-1}$ $=\kappa\frac{N!(M-N)!}{M!}$ (10)
$=2^{-(\gamma-1)N\log N-O(N)},$ (11)
for constant $\beta$ and so as $\kappa$, where we used the relation $M=c_{0}N^{\gamma}$ with a constant $c_{0}$ and $\gamma\geq 1$. By setting $\eta$ and $\xi$ small constant satisfying $\eta+\frac{\xi}{\eta}<\frac{1}{4}$, we can solve the problem in Theorem 4 up to additive imprecision $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$ using the oracle $\mathcal{O}$. Hence, assuming that the above problem is #P-hard under $\rm{BPP}^{\rm{NP}}$ reduction, we can obtain the complexity-theoretical relation $\rm{P}^{\#P}\subseteq\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$, which implies the collapse of PH if $\mathcal{S}$ with respect to shallow-depth circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ can be done in classical polynomial time. This completes the proof. ∎
VI Classical simulation hardness of shallow-depth Gaussian boson sampling
In this section, we show that our hardness results of the shallow-depth boson sampling can be generalized to the Gaussian boson sampling scheme  [2]. Our specific setup for the Gaussian boson sampling is as follows. Let the total mode number $M$ of the circuit be a power of 2, and now the input state is an $M$ product of single-mode squeezed vacuum (SMSV) state |$\ket{\text{SMSV}}^{\otimes M}$ with equal squeezing parameter $r$ and equal squeezing direction. Also, let us define the output mean photon number as an integer $N$ (i.e., $N=M\sinh^{2}r$) where $M$ and $N$ are polynomially related as $M=c_{1}N^{\gamma}$ for a constant $c_{1}$ and $\gamma\geq 1$. We define $q_{\bm{s}}(C)$ as an output probability of the Gaussian boson sampling, for an $N$ photon outcome $\bm{s}$ from an $M$ mode linear optical circuit matrix $C$ on input $M$ SMSV states. For collision-free outcome $\bm{s}$, $q_{\bm{s}}(C)$ can be expressed as [2]
$q_{\bm{s}}(C)=\left|\bra{\bm{s}}\hat{\mathcal{U}}(C)\ket{\text{SMSV}}^{\otimes M}\right|^{2}=\frac{\tanh^{N}r}{\cosh^{M}r}|\text{Haf}((CC^{T})_{\bm{s}})|^{2},$ (12)
where |$\ket{\bm{s}}$ is an $M$-mode Fock state corresponding to the outcome $\bm{s}$, $\hat{\mathcal{U}}(C)$ is a unitary operator corresponding to the circuit $C$, and $(CC^{T})_{\bm{s}}$ is an $N$ by $N$ matrix obtained by taking $s_{i}$ copies of the $i$th row and column of the matrix $CC^{T}$.
Using the above settings, we first prove the worst-case hardness of Gaussian boson sampling for a fixed outcome $\bm{s}$, with the shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
Theorem 6.
Approximating the output probability $q_{\bm{s}_{0}}(C)$ of Gaussian boson sampling to within additive error $2^{-\frac{\gamma-1}{2}N\log N-O(N)}$ for any $C$ over linear optical circuit architecture $\mathcal{B}\mathcal{B}^{*}$ is #$\rm{P}$-hard in the worst case.
Proof.
We establish a reduction from the worst-case hardness of boson sampling in Theorem 3 to the problem in Theorem 6. Let $p_{\bm{s}_{0}}(C_{0})$ be the output probability of a fixed input and output $\bm{s}_{0}$ of boson sampling in Theorem 3, for mode number $M_{0}$, photon number $N_{0}$, and the circuit $C_{0}$ in $M_{0}$ mode circuit architecture $\mathcal{B}\mathcal{B}^{*}$. In the following, we show that $p_{\bm{s}_{0}}(C_{0})$ can be efficiently reduced to the output probability $q_{\bm{s}}(C)$ of Gaussian boson sampling, for mode number $M=2M_{0}$ and mean photon number $N=2N_{0}$, with output $\bm{s}$ and circuit $C$ determined by $\bm{s}_{0}$ and $C_{0}$ each.
Our strategy is to employ the scheme in Ref. [40], which used $M_{0}$ product of equally squeezed two-mode squeezed vacuum (TMSV) state as an input state to perform $M_{0}$ mode boson sampling task. Specifically, a single TMSV state with squeezing parameter $r$ can be represented as
|$\ket{\text{TMSV}}=\frac{1}{\cosh r}\sum_{n=0}^{\infty}\tanh^{n}r\ket{n}\ket{n},$ (13)
and thus $M_{0}$ product of the TMSV state is
|$\begin{split}\ket{\text{TMSV}}^{\otimes M_{0}}&=\frac{1}{\cosh^{M_{0}}r}\left(\sum_{n=0}^{\infty}\tanh^{n}r\ket{n}_{(1)}\ket{n}_{(2)}\right)^{\otimes M_{0}}\\ &=\frac{1}{\cosh^{M_{0}}r}\sum_{n=0}^{\infty}\tanh^{n}r\sum_{\bm{s}_{n}}\ket{\bm{s}_{n}}_{(1)}\ket{\bm{s}_{n}}_{(2)},\end{split}$ (14)
where the summation of $\bm{s}_{n}$ is over all possible configurations of Fock state with a total $M_{0}$ mode and $n$ photon.
For each mode in the given $M_{0}$ mode circuit $C_{0}$, one-half of the TMSV state (i.e., subscript (2) in Eq. (14)) is input into it, and the other half of each state (i.e., subscript (1) in Eq. (14)) is sent directly to a photon counter. By setting each |$\ket{\bm{s}_{\text{in}}}$ and |$\ket{\bm{s}_{\text{out}}}$ as a total $M_{0}$ mode and total $N_{0}$ photon Fock state, the output probability can be represented as
$\left|\bra{\bm{s}_{\text{in}}}_{(1)}\bra{\bm{s}_{\text{out}}}_{(2)}\hat{\mathcal{U}}_{(2)}(C_{0})\ket{\text{TMSV}}^{\otimes M_{0}}\right|^{2}=\frac{\tanh^{2N_{0}}r}{\cosh^{2M_{0}}r}\left|\bra{\bm{s}_{\text{in}}}\hat{\mathcal{U}}(C_{0})\ket{\bm{s}_{\text{out}}}\right|^{2},$ (15)
which is the output probability of $M_{0}$ mode and $N_{0}$ photon boson sampling in circuit $C_{0}$, with an additional multiplicative factor.
Note that two $M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture can be embedded in the middle of an $M=2M_{0}$ mode $\mathcal{B}\mathcal{B}^{*}$ architecture. Accordingly, we define a circuit $C$ in $M$ mode $\mathcal{B}\mathcal{B}^{*}$ by embedding the given $M_{0}$ mode circuit $C_{0}$ in one side of $\mathcal{B}\mathcal{B}^{*}$, setting gates located right in front of the input ports as balanced beam splitters, and setting the remaining gates as identity gates; we leave in Fig. 3 an illustration of $M=16$ mode circuit $C$ for more clarity. Here, the input $M$ SMSV states with squeezing parameter $r$ combined with the balanced beam splitters at the front becomes $M_{0}$ TMSV states with squeezing parameter $r$. Therefore, our overall setup exactly follows the scheme in Ref. [40], such that the first $M_{0}$ mode is the photon counter sector to determine the input configuration of boson sampling, and the last $M_{0}$ mode is to simulate the boson sampling for the given circuit $C_{0}$.
Figure 3: Schematics of an mode number $M=16$ circuit $C$ in $\mathcal{B}\mathcal{B}^{*}$ which contains a given $M_{0}=8$ mode circuit $C_{0}$ in $\mathcal{B}\mathcal{B}^{*}$
We also define an $M$-dimensional vector $\bm{s}$ as a serial concatenation of two $\bm{s}_{0}$ vectors, so that $\bm{s}$ represents $N$ photon outcome over $M$ modes. Then the output probabilities $p_{\bm{s}_{0}}(C_{0})$ and $q_{\bm{s}}(C)$ are related as
$\begin{split}q_{\bm{s}}(C)&=\left|\bra{\bm{s}}\hat{\mathcal{U}}(C)\ket{\text{SMSV}}^{\otimes M}\right|^{2}\\ &=\left|\bra{\bm{s}_{0}}_{(1)}\bra{\bm{s}_{0}}_{(2)}\hat{\mathcal{U}}_{(2)}(C_{0})\ket{\text{TMSV}}^{\otimes M_{0}}\right|^{2}\\ &=\frac{\tanh^{2N_{0}}r}{\cosh^{2M_{0}}r}p_{\bm{s}_{0}}(C_{0}).\end{split}$ (16)
Hence, approximating the output probability $p_{\bm{s}_{0}}(C_{0})$ can be reduced to approximating the output probability $q_{\bm{s}}(C)$ of Gaussian boson sampling, with a blowup in the additive imprecision. The size of the additive imprecision blowup is
$\frac{\cosh^{2M_{0}}r}{\tanh^{2N_{0}}r}=\left(\frac{M+N}{M}\right)^{M_{0}+N_{0}}\left(\frac{M}{N}\right)^{N_{0}}=2^{\frac{\gamma-1}{2}N\log N+O(N)},$ (17)
using the relation $N=M\sinh r$ and $M\propto N^{\gamma}$. Since the allowed additive error for the worst-case hardness of $p_{\bm{s}_{0}}(C_{0})$ is $2^{-O(N)}$, the allowed additive error for the reduction is $2^{-O(N)}2^{-\frac{\gamma-1}{2}N\log N-O(N)}$ = $2^{-\frac{\gamma-1}{2}N\log N-O(N)}$. This completes the proof. ∎
Using the results of Theorem 6 and the previous proof of the average-case hardness of boson sampling in Theorem 4, it is straightforward to find the average-case hardness of Gaussian boson sampling, for randomly chosen $N$ photon outcomes $\bm{s}\sim\mathcal{G}_{M,N}$ and randomly chosen circuits $U\sim\mathcal{H}_{\mathcal{A}}$ in shallow-depth architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$.
Theorem 7.
The following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a random circuit $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ and a random outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the output probability $q_{\bm{s}}(U)$ of Gaussian boson sampling within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
Proof.
The procedure is the same as the proof of Theorem 4, namely, establishing a worst-to-average-case reduction from the problem in Theorem 6 to the problem in Theorem 7. The only different part for the Gaussian boson sampling case is the functional form of the output probability $q_{\bm{s}}(U(\theta))$ parameterized by $\theta$. Hence, we show that $q_{\bm{s}}(U(\theta))$ can also be represented as a degree $(4mN,4mN)$ rational function in $\theta$, the same degree as the boson sampling case in Lemma 3.
From Eq. (12), the output probability $q_{\bm{s}}(U(\theta))$ has the form of
[…]
where $\mu$ is along all possible perfect matching permutations over $N$ modes. From the proof of Lemma 3, using reduction to the common denominator for all of the $m=qM\log M$ gates, $\left[U(\theta)\right]_{j,k}$ can be represented as $(2m,2m)$ rational function in $\theta$ with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)$. Using this fact, one can easily check that $\prod_{j=1}^{N/2}\left[(U(\theta)U(\theta)^{T})_{\bm{s}}\right]_{\mu(2j-1),\mu(2j)}$ can be represented as $(2mN,2mN)$ rational function in $\theta$, with the common denominator $[\prod_{i=1}^{m}q_{i}(\theta)]^{N}$ which does not change with $\mu$. Therefore, the output probability can be represented as $q_{\bm{s}}(U(\theta))=\frac{P(\theta)}{Q(\theta)}$, with each $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$ and $P(\theta)$ a degree $4mN$ polynomial function in $\theta$.
Given that $q_{\bm{s}}(U(\theta))$ can be represented as a degree $(4mN,4mN)$ rational function with the same denominator $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$ from the boson sampling case in Lemma 3, we can repeat all the steps identically to the proof of Theorem 4 and obtain the same result. ∎
VII Extension of hardness results for lossy environments
[…]
To proceed, we start with a brief review of the results presented by Ref. [41], which shows the hardness of simulating noisy quantum circuits. Specifically, one can simulate a noiseless circuit using a larger noisy circuit up to the desired imprecision, by establishing error-detecting code in the noisy circuit and post-selecting null syndrome measurements. Therefore, given the probability to post-select the no-error syndromes, one can approximate the output probability of the noiseless circuit from the output probability of the noisy circuit. Based on this argument, Ref. [21] demonstrates the average-case hardness of approximating output probabilities of noisy quantum circuits, under some plausible assumptions of the noise model. This result gives evidence of the approximate simulation hardness of noisy quantum circuits, within total variation distance error.
[…]
For a more detailed analysis, we set the loss model as follows. Let the photon loss model $\mathcal{N}$ be local and stochastic. Specifically, $\mathcal{N}$ is a set of loss channels $\{\mathcal{N}_{i}\}_{i=1}^{l}$, such that after each unitary gate is applied, loss channel $\mathcal{N}_{i}$ acts on each mode participated in the unitary gate. Hence, the number of loss channels is $l=O(m)$ for gate number $m$ in a given circuit architecture. We can decompose each noise channel $\mathcal{N}_{i}$ as follows:
$\mathcal{N}_{i}=(1-\rho_{i})\mathcal{I}+\rho_{i}\mathcal{E}_{i},$ (19)
where $\mathcal{I}$ is identity, $\mathcal{E}_{i}$ is an CPTP map representing photon loss, and $\rho_{i}$ is a loss rate for each channel satisfying $\rho_{i}\leq\rho$ for a constant $\rho$. The validity of such modeling for photon loss channel is represented in [12, 3].
[…]
$p_{\bm{s}}(C)=\frac{p_{\bm{s}}(C,\mathcal{N})}{\Pr[\text{`no loss event'}]}.$ (20)
From Eq. (19), the probability of ‘no loss event’ is $\prod_{i=1}^{l}(1-\rho_{i})$, which can be efficiently calculated. This implies that approximating $p_{\bm{s}}(C,\mathcal{N})$ can be reduced from approximating $p_{\bm{s}}(C)$, with at most $\Pr[\text{`no loss event'}]^{-1}=\prod_{i=1}^{l}(1-\rho_{i})^{-1}\leq(1-\rho)^{-l}=2^{O(\rho m)}$ blowup in the additive imprecision.
Given Eq. (20), we can repeat the same steps from the previous hardness arguments, for the lossy shallow-depth boson sampling; the only difference is the imprecision blowup by $2^{O(\rho m)}$. For our shallow-depth architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$, the gate number $m$ is $qM\log M$, so the size of imprecision blowup is $2^{O(N^{\gamma}\log N)}$ in our case. Such imprecision blowup does not affect the allowed additive accuracy $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$ for our average-case hardness result. Based on the arguments so far, the following corollary is straightforward.
Corollary 1.
Suppose we have the photon loss model $\mathcal{N}$ with each loss rate $\rho_{i}\leq\rho$ for a constant $\rho$. Then the following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a random circuit $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ and a random outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the lossy output probability $p_{\bm{s}}(U,\mathcal{N})$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
We remark that for our noise model, the imprecision blowup grows exponentially with the gate number $m$. Therefore, shallow-depth circuits can be more advantageous in this perspective, since they are likely to have less gate number and thus have small imprecision blowup. For example, the current hardness results are based on $M$ by $N$ submatrices of $M$-dimensional Haar random unitaries, and the implementation of such matrices requires gate number $m=\Omega(N^{\gamma+1})$. This arouses the imprecision blowup at least $2^{O(N^{\gamma+1})}$, which restricts the allowed additive error for the average-case hardness at most $2^{-O(N^{\gamma+1})}$.
VIII Concluding remarks
Here we provide a few remarks about our overall results and related open questions.
1. Our result demonstrates the average-case hardness for additive imprecision $2^{-O(N^{\gamma+1}(\log N)^{2})}$. Indeed, there still remains a gap to the desired additive imprecision for the simulation hardness $2^{-(\gamma-1)N\log N-O(N)}$ in Theorem 5. Hence, closing this gap would be an ultimate challenge to the full achievement of classical intractability; more advanced proof techniques are required to reduce this gap. Here, one can take the following approach: finite-size numerical experiments suggest that the output distributions of local random circuits in the butterfly circuit architecture (Definition 2) are close enough to those of global Haar random circuits [37]. Accordingly, if one can analytically prove that the distance between those output distributions is close enough, we can directly obtain a better imprecision level $2^{-O(N\log N)}$ by results in [21, 3, 22], which employed the global Haar random circuits.
Another possible approach for reducing the imprecision gap is to perturb a random circuit matrix in a different way from the Cayley transform (Definition 6), i.e., as depicted in [22]. Specifically, instead of perturbing each random gate, one can perturb a submatrix $X$ of our random circuit matrix $U$ with a worst-case matrix $A$ as $X(\theta)=(1-\theta)X+\theta A$ for $\theta\in[0,1]$. Here, a degree of polynomial $|\text{Per}(X(\theta))|^{2}$ is $2N$, which is lower than ours derived by the Cayley transform. Therefore, if one can prove that $X(\theta)$ is distributed similarly to $X$ for small $\theta$, we expect that we can also obtain a better imprecision level by using the same interpolation method. However, the above approach requires one to figure out a global circuit distribution generated by the convolution of local circuit distributions. Although we believe that this problem can be resolved using techniques from random matrix theories, we have not yet developed a complete analysis. Hence, we leave it as an open question.
2. Another important challenge that should be addressed is to find the classical simulation hardness of noisy boson sampling, for general types of physical noise beyond the photon loss noise model we have dealt with so far. To do so, as described in [41, 21, 3], employing the threshold theorem would be a viable choice for this goal. Specifically, the threshold theorem for general types of noise in boson sampling setups has to be developed. This requires an efficient error detection code for general types of error using linear optical elements, for any multi-mode Fock state or Gaussian state input. However, to the best of our knowledge, such an error detection code does not exist. Hence, constructing this error detection code would be a crucial step toward the hardness of noisy boson sampling, which will contribute to a more noise-tolerant demonstration of quantum advantage with boson sampling. We leave this problem as another open question.
Acknowledgement
We thank Bill Fefferman for insightful discussions. The authors acknowledge support from the National Research Foundation of Korea (NRF) grants funded by the Korean government (Grant Nos. NRF-2023R1A2C1006115, NRF-2022M3K4A1097117, and NRF-2022M3E4A1076099) via the Institute of Applied Physics at Seoul National University, the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2021-0-01059 and IITP-20232020-0-01606). B.G. was also supported by the education and training program of the Quantum Information Research Support Center, funded through the National Research Foundation of Korea (NRF) by the Ministry of Science and ICT (MSIT) of the Korean government (No.2021M3H3A1036573).
Appendix A Previous foundations: Average-case hardness of boson sampling
In this appendix, we argue the existing proof technique employed for the simulation hardness of boson sampling, specifically in the context of the approximate simulation within total variation distance error [1]. The current state-of-the-art proof technique for the hardness of sampling problems like boson sampling essentially builds upon Stockmeyer’s algorithm about approximate counting [38]. Specifically, given a classical sampler that outputs a sample from a given output distribution, Stockmeyer’s algorithm enables one to multiplicatively estimate a fixed output probability of the sampler, within complexity class $\rm{BPP}^{\rm{NP}}$.
Now suppose there exists an approximate classical sampler capable of simulating ideal boson sampling up to total variation distance error, as in Definition 8. This approximate sampler can have a large additive error for a fixed output probability, but have a comparably small additive error for most of the output probabilities due to Markov’s inequality. Then, Stockmeyer’s algorithm, combined with the approximate sampler, can well approximate the ideal output probability of boson sampling within a certain additive error, with a high probability over randomly chosen outcomes (See Lemma 4 for more details). For convenience, let us refer to this computational task as an average-case approximation problem of boson sampling. If the complexity of the average-case approximation problem is outside the Polynomial Hierarchy (PH), it implies the collapse of PH, since the complexity of Stockmeyer’s algorithm is indeed inside the finite level of PH.
Here, average-case hardness comes into the proof of the classical simulation hardness argument, which means that approximating the ideal output probability of boson sampling with high probability over randomly chosen outcomes is #P-hard. More precisely, if the average-case hardness holds up to the imprecision level of the average-case approximation problem, this comes down to the classical simulation hardness of the approximate sampling unless PH collapses, by the complexity-theoretical foundation $\rm{PH}\subseteq\rm{P}^{\#P}$ [39].
Moreover, by choosing random circuit instances that have symmetry over the outcomes, one can reduce the average-case instances for the hardness from outcome instances to circuit instances, which is called the hiding property. For the boson sampling case, global Haar random unitary (i.e., unitary matrix drawn from Haar measure on U(M), for mode number M) satisfies this condition. In detail, instead of randomly choosing the outcome, we can fix the outcome by applying a random permutation to the global Haar random unitary distribution, which is still Haar distributed from its symmetric property. This hiding property plays an important role in the current proofs of the average-case hardness, as it enables one to establish worst-to-average-case reduction. Specifically, as the output probability of boson sampling can be written as a low-degree polynomial of input circuit (matrix) values, it allows one to infer the value of a worst-case instance from the output probability of many average-case circuit instances. Hence, the average-case hardness argument for boson sampling is typically used in this context, i.e., average-case hardness over random circuit instances, for a fixed outcome [1, 3, 21].
Accordingly, the crucial step for the classical simulation hardness of approximate sampling is to prove the average-case hardness for the desired imprecision level. While there have been many impressive results about the average-case hardness of boson sampling [1, 3, 21, 22], the average-case hardness for the desired imprecision level is not yet fully demonstrated. Still, there exists a gap between the imprecision level of average-case hardness in the strongest existing results and the imprecision level of average-case approximation problem. Hence, closing this imprecision gap remains the ultimate challenge for the fully theoretically guaranteed computational advantage of approximate boson sampling.
Appendix B Proof of Theorem 3
The proof of Theorem 3 can be greatly simplified by introducing the following two Lemmas.
Lemma 5 (Brod [25], revised).
For an arbitrary given poly-sized $n$-qubit quantum circuit $Q$, there exists a constant depth linear optical circuit $C$ such that for $M\geq 2N$ and $N=\text{poly}(n)$,
[…]
where $c_{Q}$ is a $Q$ dependent constant which can be efficiently computed, $\hat{\mathcal{U}}(C)$ is a unitary operator corresponding to the circuit $C$, and |$\ket{I}$ is an $M$-mode Fock state composed of $N$ single photon states and vacuum states for the rest modes.
Proof.
We revise the results by [25], for a more rigorous analysis of the allowed additive imprecision level for the worst-case hardness of shallow-depth boson sampling. Ref. [25] proposed that certain 4-depth linear-optical circuits with post-selection can simulate universal quantum computing. Specifically, to simulate any poly-sized quantum circuit $Q$ on $n$ qubits, there exists a measurement-based quantum computation (MBQC) scheme using constant depth brickwork graph state of maximally poly($n$) qubits [42, 43, 44]. The corresponding scheme can also be implemented in a linear optical system via KLM scheme [45] with post-selection, using $N=\text{poly}(n)$ number of single photon states over $M\geq 2N$ modes (requirements for the dual-rail encoding), which can simulate the quantum circuit composed of $O(N)$ number of gates.
Therefore, given the circuit in [25] and an appropriate dual-rail encoded state |$\ket{I}$, the output probability of any quantum circuit $Q$ can be represented as Eq. (21), where $c_{Q}$ denotes the product of post-selection probabilities for gate implementations. To compute $c_{Q}$, we need to figure out the required number of gates to implement the circuit $Q$ and their probabilities to be post-selected. More precisely, from [25], post-selection occurs for two cases: $(i)$ for the CZ gate to implement the brickwork graph state, and $(ii)$ for the gate set {CX, T, H} (i.e., universal set of gates) which can be implemented by measurement of the graph state.
Hence, $c_{Q}$ can be expressed as
$c_{Q}$ $=\prod_{k\in\{\text{CZ, CX, T, H}\}}p_{k}^{\Gamma_{k}},$ (22)
where $p_{k}$ denotes post-selection probability of $k$ gate (e.g., $p_{\text{CZ}}$ is $2/27$ in [46]), and $\Gamma_{k}$ denotes the number of $k$ gate to implement the circuit $Q$. By counting the number of each gate to implement the circuit $Q$, $c_{Q}$ can be computed efficiently. ∎
Lemma 6 (Kondo et al [34]).
It is #$\rm{P}$-hard to compute $|\bra{0}^{\otimes n}Q\ket{0}^{\otimes n}|^{2}$ for an arbitrary given quantum circuit $Q$ within the additive error less than $2^{-2n}$.
Combining the above results, now we prove the worst-case hardness of output probability approximation of boson sampling in the shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$, for fixed input and output corresponding to the |$\ket{I}$ in Lemma 5. By Lemma 6, approximating output probabilities of worst-case $n$-qubit BQP circuits within additive error $2^{-2n}$ is #P-hard. Also, one can easily check that the constant depth linear-optical circuit proposed by [25] can be efficiently embedded in the architecture $\mathcal{B}\mathcal{B}^{*}$. Hence, by Lemma 5, approximating $p_{\bm{s}}(C)$ of any $C$ over $\mathcal{B}\mathcal{B}^{*}$, for both input and output $\bm{s}$ corresponding to the |$\ket{I}$, can be reduced from approximating output probabilities of any BQP circuits, with $c_{Q}^{-1}$ blowup in the additive imprecision. Since the post-selection occurs $O(N)$ times to implement $n$-qubit BQP circuit by Lemma 5, $c_{Q}$ has its amplitude $c_{Q}=2^{-O(N)}$, and thus the allowed additive error for the reduction is $2^{-(2n+O(N))}=2^{-O(N)}$.
Appendix C Proof of Theorem 4
In this proof, we establish the worst-to-average-case reduction, from the problem in Theorem 3 to the problem in Theorem 4. Let $\mathcal{O}$ be an oracle that solves the problem in Theorem 4, i.e., on input $\bm{s}\sim\mathcal{G}_{M,N}$ and $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ for fixed $q\geq 2$, the oracle outputs $p_{\bm{s}}(U)$ within additive error $\epsilon$ for high constant probability over $U$ and $\bm{s}$. Let $C_{0}$ be the worst-case circuit in $(\mathcal{B}\mathcal{B}^{*})^{q_{0}}$ with $q_{0}\geq 1$, and $\bm{s}_{0}$ the fixed collision-free output given in Theorem 3. In the following, we show that approximating $p_{\bm{s}_{0}}(C_{0})$ to within additive error $2^{-O(N)}$ (i.e., Theorem 3) is in $\text{BPP}^{\text{NP}^{\mathcal{O}}}$, which implies that the average-case approximation of $p_{\bm{s}}(U)$ to within $\epsilon$ is #P-hard under $\text{BPP}^{\text{NP}}$ reduction.
[…]
For the reduction from the worst-case circuit $C_{\bm{P}}$ to the average-case circuits, we sample random circuit $U(\theta)$ in $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $q=q_{0}+1$, by sampling independently distributed local Haar random gate $\{H_{i}\}_{i=1}^{m}$ for gate number $m=qM\log M$, perturbing them by the Cayley transform parameterized by $\theta$ and multiplying the worst-case circuit gates from $C_{\bm{P}}$ as in Definition 7. Then, $U(\theta)$ follows the distribution $\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}}$, with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$, and $U(1)=C_{\bm{P}}$.
Given randomly chosen outcome $\bm{s}_{\bm{P}}\sim\mathcal{G}_{M,N}$ and circuit $U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}}$, we input them in the oracle $\mathcal{O}$. For at least $1-\eta$ over $\bm{s}_{\bm{P}}$, the failure probability of $\mathcal{O}$ is at most
$\Pr_{U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}}}[|\mathcal{O}(\bm{s}_{\bm{P}},U(\theta))-p_{\bm{s}_{\bm{P}}}(U(\theta))|>\epsilon]<\delta+D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}},\mathcal{H}_{\mathcal{A}}),$ (23)
where $D_{\rm{TV}}$ denotes total variation distance. This is evident as we can interpret the total variation distance as the supremum over events of the difference in probabilities of those events (Viz., circuits corresponding to the failure) [32]. By Lemma 2, $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}},\mathcal{H}_{\mathcal{A}})$ is $O(m\theta)$. By setting $0\leq\theta\leq\Delta$ with $\Delta=O(m^{-1})$, we can upper bound $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}},\mathcal{H}_{\mathcal{A}})$ by an arbitrarily small constant.
By Lemma 3, $p_{\bm{s}_{\bm{P}}}(U(\theta))$ is a $(4mN,4mN)$ degree rational function $\frac{P(\theta)}{Q(\theta)}$, where the denominator is given as $Q(\theta)=\left[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}\right]^{N}$. We note that $Q(\theta)$ can be computed in $\Theta(m)$ time, as it only depends on the constant number of eigenvalues of local gate matrices (i.e., $\phi_{j}$ values in Eq. (4) for each local gate matrix $H_{i}$). Also, given that $\theta\leq\Delta=O(m^{-1})$, $Q(\theta)$ is very close to the unity, as $Q(\theta)\geq 1$ and
$Q(\theta)$ $=\left[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}\right]^{N}$
$=\left[\prod_{i=1}^{m}\prod_{j=1}^{2}|(1+i\theta e^{i\frac{\phi_{i,j}}{2}}\sin{\frac{\phi_{i,j}}{2}})|^{2}\right]^{N}$ (24)
$\leq\left[\prod_{i=1}^{m}\prod_{j=1}^{2}(1+\theta^{2})\right]^{N}$
$\leq(1+O(m^{-2}))^{2mN}$
$=1+O(Nm^{-1}),$
where $\phi_{i,j}$ denotes the phase of the $j$th eigenvalue of the $i$th gate.
Therefore, $p_{\bm{s}_{\bm{P}}}(U(\theta))$ is very close to the degree $d=4mN$ polynomial $P(\theta)$ in $\theta\in[0,\Delta]$, which allows us to use polynomial interpolation technique for $P(\theta)$. Specifically, we obtain estimations of $P(\theta)$ for different values of $\theta\in[0,\Delta]$ by querying the oracle $\mathcal{O}$, use polynomial interpolation for given $P(\theta)$ values to estimate $P(1)$, and infer the value $p_{\bm{s}_{\bm{P}}}(U(1))=p_{\bm{s}_{\bm{P}}}(C_{\bm{P}})$ by multiplying $Q(1)^{-1}$. However, $Q(1)^{-1}$ becomes arbitrarily large for the case that even a single $\phi_{i,j}$ in Eq. (C) is near ±$\pm\pi$, which will arbitrarily enlarge the imprecision of the approximation of $p_{\bm{s}_{\bm{P}}}(U(1))$. To avoid this issue, we employ the strategy from Ref. [21], which only considers the case that all $\phi_{i,j}$ values of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ are in $[-\pi+\zeta,\pi-\zeta]$, and regards the other case as failure. This happens with probability at least $1-O(m\zeta)$ over the random circuit instances. By setting $\zeta=O(m^{-1})$, we can make $O(m\zeta)$ arbitrarily small constant, and as a result, we can upper bound $Q(1)^{-1}$ (see Eq. (28) below) with high probability over random circuit instances.
Now the problem reduces to approximating degree $d=4mN$ polynomial $P(\theta)$ with the value $\mathcal{O}(\bm{s}_{\bm{P}},U(\theta))Q(\theta)$ in $\theta\in[0,\Delta]$ within additive error smaller than $\epsilon(1+O(Nm^{-1}))\approx\epsilon$; such approximations will later be used for the estimation of the value $P(1)$ via polynomial interpolation technique. The failure of $\mathcal{O}$ depends on the outcome $\bm{s}_{\bm{P}}\sim\mathcal{G}_{M,N}$ whose failure probability is at most $\eta$, and the circuit $U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{\bm{P}}}$ whose failure probability is at most $\delta+O(m\Delta)$ from Eq. (23). Also, the probability that at least one $\phi_{i,j}$ of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ is outside of the regime $[-\pi+\zeta,\pi-\zeta]$ is at most $O(m\zeta)$. Putting everything together and applying a simple union bound, the total failure probability of the approximation of $P(\theta)$ is at most
$\begin{split}\Pr[|\mathcal{O}(\bm{s}_{\bm{P}},U(\theta))Q(\theta)-P(\theta)|>\epsilon]&<\eta+\delta+O(m\Delta)+O(m\zeta)\\ &\leq\delta^{\prime},\end{split}$ (25)
where $\delta^{\prime}$ is an upper bound of $\eta+\delta+O(m\Delta)+O(m\zeta)$, and given $\eta+\delta<\frac{1}{4}$, we can make $\delta^{\prime}<\frac{1}{4}$ by setting $O(m\Delta)$ and $O(m\zeta)$ arbitrary small constants.
Let $\{\theta_{i}\}_{i=1}^{O(d^{2})}$ be the set of equally spaced points in the interval $[0,\Delta]$. For each $\theta_{i}$, we obtain the unitary matrix $U(\theta_{i})$ using the same random gate $\{H_{i}\}_{i=1}^{m}$ and worst-case circuit $C_{\bm{P}}$. Let $y_{i}=\mathcal{O}(\bm{s}_{\bm{P}},U(\theta_{i}))Q(\theta_{i})$. By Eq. (25), each set of points $(\theta_{i},y_{i})$ satisfies
$\Pr[|y_{i}-P(\theta_{i})|>\epsilon]\leq\delta^{\prime}<\frac{1}{4}.$ (26)
By using the interpolation algorithm introduced in Theorem 8, we can obtain the additive approximation of $P(1)$ as $\tilde{p}$ with an access to NP oracle, such that
$\Pr[|\tilde{p}-P(1)|>\epsilon^{\prime}]<\frac{1}{3},$ (27)
where $\epsilon^{\prime}=\epsilon e^{-d\log{\Delta}}=\epsilon 2^{O(N^{\gamma+1}(\log N)^{2})}$ using $d=4mN$ and $m=qM\log M$. Note that the failure probability in Eq. (27) can be arbitrarily reduced by taking a polynomial number of trials, and thus we can obtain the estimated value $P(1)$ within additive error $\epsilon^{\prime}$ with arbitrarily high probability.
From the estimated value $P(1)$, we can infer the worst-case output probability value $p_{\bm{s}_{\bm{P}}}(U(1))=P(1)/Q(1)$. As the value of $Q(1)$ depends on the values $\phi_{i,j}$ in Eq. (C), the $\phi_{i,j}$ independent lower bound of $Q(1)$ is required to set an upper bound of the additive imprecision of $p_{\bm{s}_{\bm{P}}}(U(1))$. Since we only consider the case that all of $\phi_{i,j}$ values are in $[-\pi+\zeta,\pi-\zeta]$ with $\zeta=O(m^{-1})$ for all randomly chosen gates $\{H_{i}\}_{i=1}^{m}$, we have
$\begin{split}Q(1)&=\left[\prod_{i=1}^{m}\prod_{j=1}^{2}|(1+ie^{i\frac{\phi_{i,j}}{2}}\sin{\frac{\phi_{i,j}}{2}})|^{2}\right]^{N}\\ &=\left[\prod_{i=1}^{m}\prod_{j=1}^{2}\left(1-\sin^{2}{\frac{\phi_{i,j}}{2}}\right)\right]^{N}\\ &\geq\left[\prod_{i=1}^{m}\prod_{j=1}^{2}\left(1-\sin^{2}{\frac{\pi-\zeta}{2}}\right)\right]^{N}\\ &=\left(O(m^{-2})\right)^{2mN}\\ &=2^{2mN\log O(m^{-2})}.\end{split}$ (28)
Therefore, the total additive error for estimating $p_{\bm{s}_{\bm{P}}}(U(1))$ is bounded by $\epsilon^{\prime}2^{-2mN\log O(m^{-2})}=\epsilon 2^{O(N^{\gamma+1}(\log N)^{2})}$. By setting $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}2^{-O(N)}=2^{-O(N^{\gamma+1}(\log N)^{2})}$, we can estimate the worst-case output probability value $p_{\bm{s}_{\bm{P}}}(U(1))=p_{\bm{s}_{\bm{P}}}(C_{\bm{P}})=p_{\bm{s}_{0}}(C_{0})$ within additive error $2^{-O(N)}$, and the whole reduction process is in $\text{BPP}^{\text{NP}}$. This completes the proof.
For the polynomial interpolation, we employ the Robust Berlekamp-Welch algorithm recently proposed in Ref. [21].
Theorem 8 (Robust Berlekamp-Welch [21]).
Let $P(x)$ be a degree $d$ polynomial in $x$. Suppose there is a set of points $D=\{(x_{i},y_{i})\}$ such that $|D|=O(d^{2})$ and $\{x_{i}\}$ is equally spaced in the interval $[0,\Delta]$. Suppose also that each points $(x_{i},y_{i})$ satisfies
$\Pr[|y_{i}-P(x_{i})|\geq\epsilon]\leq\delta,$ (29)
with $\delta<\frac{1}{4}$. Then there exists a $\rm{P}^{\rm{NP}}$ algorithm that takes input $D$ and outputs $\tilde{p}$ such that
$|\tilde{p}-P(1)|\leq\epsilon e^{-d\log\Delta},$ (30)
with success probability at least $\frac{2}{3}$.
Appendix D Proof of Lemma 4
Let $\bar{p}_{\bm{s}}(C)$ be the output probability distribution from the approximate sampler $\mathcal{S}$ with the given linear optical circuit $C$. Also, let $\mathcal{C}_{M,N}$ be the set of collision-free outcomes of boson sampling, for mode number $M$ and photon number $N$. Then $\bar{p}_{\bm{s}}(C)$ satisfies
$\operatorname*{\mathbb{E}}_{\bm{s}\sim\mathcal{G}_{M,N}}\left[|\bar{p}_{\bm{s}}(C)-p_{\bm{s}}(C)|\right]=\frac{1}{\binom{M}{N}}\sum_{\bm{s}\in\mathcal{C}_{M,N}}|\bar{p}_{\bm{s}}(C)-p_{\bm{s}}(C)|\leq\frac{2\beta}{\binom{M}{N}}.$ (31)
Using Eq. (31) and Markov’s inequality, $\bar{p}_{\bm{s}}(C)$ satisfies
$\Pr_{\bm{s}\sim\mathcal{G}_{M,N}}\left[|\bar{p}_{\bm{s}}(C)-p_{\bm{s}}(C)|\geq\frac{\beta k}{\binom{M}{N}}\right]$ $\leq\frac{2}{k}$ (32)
for all $k>2$. Also, using Stockmeyer’s algorithm [38] whose complexity is in $\rm{BPP}^{\rm{NP}}$, obtaining the estimate $\tilde{p}_{\bm{s}}(C)$ of $\bar{p}_{\bm{s}}(C)$ satisfying
$\Pr\left[|\tilde{p}_{\bm{s}}(C)-\bar{p}_{\bm{s}}(C)|\geq\alpha\bar{p}_{\bm{s}}(C)\right]\leq\frac{1}{2^{N}},$ (33)
in polynomial time in $N$ and $\alpha^{-1}$ is in $\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Using $\operatorname*{\mathbb{E}}_{\bm{s}\sim\mathcal{G}_{M,N}}[\bar{p}_{\bm{s}}(C)]=\binom{M}{N}^{-1}\sum_{\bm{s}\in\mathcal{C}_{M,N}}\bar{p}_{\bm{s}}(C)\leq\binom{M}{N}^{-1}$,
$\Pr\left[|\tilde{p}_{\bm{s}}(C)-\bar{p}_{\bm{s}}(C)|\geq\frac{\alpha l}{\binom{M}{N}}\right]$ $\leq\Pr\left[\bar{p}_{\bm{s}}(C)\geq\frac{l}{\binom{M}{N}}\right]+\Pr\left[|\tilde{p}_{\bm{s}}(C)-\bar{p}_{\bm{s}}(C)|\geq\alpha\bar{p}_{\bm{s}}(C)\right]$ (34)
$\leq\frac{1}{l}+\frac{1}{2^{N}},$ (35)
for all $l>1$. Putting all together, by applying a triangular inequality, finding an average-case approximation $\tilde{p}_{\bm{s}}(C)$ of $p_{\bm{s}}(C)$ satisfying
$\Pr\left[|\tilde{p}_{\bm{s}}(C)-p_{\bm{s}}(C)|\geq\frac{\beta k+\alpha l}{\binom{M}{N}}\right]$ $\leq\frac{2}{k}+\frac{1}{l}+\frac{1}{2^{N}}$ (36)
is in $\rm{BPP}^{\rm{NP}^{\mathcal{S}}}$. Let $\kappa$ and $\xi$ be fixed error parameters such that $k/2=l=3/\xi$ and $\beta=\kappa\xi/{12}=\alpha/2$. As $\beta k+\alpha l=\kappa$ and $\frac{2}{k}+\frac{1}{l}+\frac{1}{2^{N}}=\frac{2}{3}\xi+\frac{1}{2^{N}}\leq\xi$, we finally obtain the Eq. (9).